Direct Square Feet of
Manufacturing Number of Floor Space Manufacturing Number of
Department Labor-Hours Employees Occupied Labor-Hours Requisitions
Building and grounds 0 0 0 0 0
Personnel* 0 0 8,400 0 0
General plant administration 0 42 12,600 0 0
Cafeteria: operating loss 0 14 5,600 1,000 0
Storeroom 0 28 39,200 2,000 0
Machining 6,000 70 56,000 6,000 6,000
Assembly 6,000 126 128,200 6,600 2,000
Total 12,000 280 250,000 15,600 8,000
*Basis used is number of employees
Support departments:
Building and grounds $35,000
Personnel 1,064
General plant administration 21,300
Cafeteria: Operating loss 3,204
Storeroom 3,388 $63,956
Operating departments:
Machining $33,000
Assembly 48,900 81,900
Total for support and operating departments $145,856

Using the​ step-down method, allocate​ support-department costs. Allocate the costs of the support departments in the order given in this problem. ​(Round all indirect allocation rates to the nearest cent and all other computations to the nearest whole dollar. Use parentheses or a minus sign when decreasing departments by allocating​ costs.)
Answer #1

Buildings and grounds

personnel

General plant administration

Cafeteria: Operating loss

Storeroom

Machining

Assembly

Step-down method

35000

1064

21300

3204

3388

33000

48900

1. buildings and ground (35000/250000) = 0. 14/square feet

(35000)

1176 (8400*0.14)

1764 (12600*0.14)

784 (5600*0.14)

5488 (39200*0.14)

7840 (56000*0.0.14)

17948 (128200*0.14)

2. personnel (2240/280) = 8/employee

(2240)

336 (42*8)

112 (14*8)

224 (28*8)

560 (70*8)

1008 (126*8)

3. General plant administration (23400/15600) = 1.50/MLH

(23400)

1500 (1000*1.50)

3000 (2000*1.50)

9000 (6000*1.5)

9900 (6600*1.5)

4. Cafeteria: Operating loss (5600/124) = 25 /employee

(5600)

700 (28*25)

1750 (70*25)

3150 (126*25)

5. Storeroom (12800/8000)= 1.6/requisition

(12800)

9600 (6000*1.6)

3200 (2000*1.6)

Cost allocated to operating departments

$61750

$84106